# Learning Autonomous Code Integration for Math Language Models

### v2

@@ Abstract
Recent advances in mathematical problem-solving with language models (LMs) integrate chain-of-thought (CoT) reasoning and code execution to harness their complementary strengths. However, existing hybrid frameworks exhibit a critical limitation: they depend on externally dictated instructions or rigid code-integration templates, lacking metacognitive awareness—the capacity to dynamically evaluate intrinsic capabilities and autonomously determine when and how to integrate tools. This rigidity motivates our study of autonomous code integration, enabling models to adapt tool-usage strategies as their reasoning abilities evolve during training.
While reinforcement learning (RL) shows promise for boosting LLM reasoning at scale (e.g., DeepSeek-R1), we demonstrate its inefficiency in learning autonomous code integration due to inadequate exploration of the vast combinatorial space of CoT-code interleaving patterns. To address this challenge, we propose a novel Expectation-Maximization (EM) framework that synergizes structured exploration (E-step) with off-policy RL optimization (M-step), creating a self-reinforcing cycle between metacognitive tool-use decisions and evolving capabilities. Experiments reveal our method achieves superior results through improved exploration. Notably, our 7B model improves over 11% on MATH500 and 9.4% on AIME without o1-like CoT. Code, models and data are released via an anonymous repository.
Learning Autonomous Code Integration for Math Language Models
Haozhe Wang†‡, Long Li†, Chao Qu†, Fengming Zhu‡, Weidi Xu†, Wei Chu†, Fangzhen Lin‡ jasper.whz@outlook.com INF Technology†, Hong Kong University of Science and Technology‡
@@ 4 Experiments
Our implementation uses rollouts per query (temperature=1.0, top-p=0.9). Training completes in about 10 hours on A100 (80GB) GPUs across three epochs of 7K queries. We release code, models and data via an anonymous repository.
@@ 5 Related Work and Discussion
Conclusion. Existing tool-integrated math language models lack the metacognitive capacity to effectively determine code integration, hindering their ability to fully realize the synergistic benefits of tool integration and CoT. To address this critical gap, we propose a novel EM-based framework that combines guided exploration with policy optimization. Our experiments demonstrate the limitations of standard SFT and RL in efficiently exploring the combinatorial space of code-integrated trajectories and highlight the superior training efficiency and performance of our approach.
Limitations. The scope of our work is primarily focused on mathematical problem-solving. While we observe promising results on challenging benchmarks like MATH500, the generalizability of our approach to other domains requiring the metacognitive capacity of tool integration and CoT, such as scientific reasoning or code generation for general-purpose tasks, remains to be explored. Future work should investigate the effectiveness of our framework across a wider range of tasks and domains.

### v3

# To Code or not to Code? Adaptive Tool Integration for Math Language Models via Expectation-Maximization

@@ Abstract
Recent advances in mathematical problem-solving with language models (LMs) integrate chain-of-thought (CoT) reasoning and code execution to harness their complementary strengths. However, existing hybrid frameworks exhibit a critical limitation: they depend on externally dictated instructions or rigid code-integration templates, lacking metacognitive awareness – the capacity to dynamically evaluate intrinsic capabilities and autonomously determine when and how to integrate tools. This rigidity motivates our study of autonomous code integration, enabling models to adapt tool-usage strategies as their reasoning abilities evolve during training.
While reinforcement learning (RL) shows promise for boosting LLM reasoning at scale (e.g., DeepSeek-R1), we demonstrate its inefficiency in learning autonomous code integration due to inadequate exploration of the vast combinatorial space of CoT-code interleaving patterns. To address this challenge, we propose a novel Expectation-Maximization (EM) framework that synergizes structured exploration (E-step) with off-policy RL optimization (M-step), creating a self-reinforcing cycle between metacognitive tool-use decisions and evolving capabilities. Experiments reveal our method achieves superior results through improved exploration. Notably, our 7B model improves over 11% on MATH500 and 9.4% on AIME without o1-like CoT. Code, models and data are released on https://github.com/HaozheH3/AutoCode.
To Code or not to Code? Adaptive Tool Integration for Math Language Models via Expectation-Maximization
Haozhe Wang1 2††thanks: Equal contribution.††thanks: Corresponding author., Long Li 111footnotemark: 1, Chao Qu 1, Fengming Zhu2, Weidi Xu1, Wei Chu 1, Fangzhen Lin222footnotemark: 2 jasper.whz@outlook.com INF Technology1, The Hong Kong University of Science and Technology2
@@ 4 Experiments
Our implementation uses rollouts per query (temperature=1.0, top-p=0.9). Training completes in about 10 hours on A100 (80GB) GPUs across three epochs of 7K queries. We list the collected public SFT data in Tab. 1.
@@ 5 Related Work and Discussion
6 Conclusion
Existing tool-integrated math language models lack the metacognitive capacity to effectively determine code integration, hindering their ability to fully realize the synergistic benefits of tool integration and CoT. To address this critical gap, we propose a novel EM-based framework that combines guided exploration with policy optimization. Our experiments demonstrate the limitations of standard SFT and RL in efficiently exploring the combinatorial space of code-integrated trajectories and highlight the superior training efficiency and performance of our approach.
7 Limitations
The scope of our work is primarily focused on mathematical problem-solving. While we observe promising results on challenging benchmarks like MATH500, the generalizability of our approach to other domains requiring the metacognitive capacity of tool integration and CoT, such as scientific reasoning or code generation for general-purpose tasks, remains to be explored. Future work should investigate the effectiveness of our framework across a wider range of tasks and domains.
Appendix A Appendix
A.1 Derivation of the EM
We first model RL as a maximum likelihood problem. Specifically, we consider ‘maximizing expected return’ as ‘maximizing the likelihood of observing a correct response‘, this is equivalent to maximizing the following log-likelihood,
where we factorize the language model as .
Since the variable is unobservable for lack of reliable supervision, we resort to the EM framework. Specifically, we treat as a hidden variable, and introduce a proposal distribution , which represents the belief of given the fixed LLM and the data evidence. We have the following derivations,
where denotes the likelihood of generating correct responses given following the solution-generation policy .
In the E-step, we are essentially minimizing a KL divergence,
where the minimizer is . Intuitively, the belief over the methodology following a posterior update: it is based on the prior of the current model and the likelihood of data . The optimal methodology-selection strategy assigns higher probability to method if following the current LLM it has higher likelihood to generate correct responses or higher prior over it.
To compute the optimal strategy, we define as an energy-based distribution, where the negative energy function evaluates the expected return induced by the current solution-generation policy: , is a temperature controlling the sharpness of the distribution. Then the minimizer can be computed by enumerating over .
In the M-step, we optimize
which maximizes the expected return while imitating the optimal strategy.
A.2 Distribution of Multi-round
We show the distribution of multi-round responses of the model AutoCode4Math-Deepseek in Tab. 3.
| Model | Dataset | Queries | Round1 | Round2 | Round3 |
| Deepseek-math-AutoCode | MATH | 5000 | 1252 | 3719 | 29 |
| GSM8k | 1319 | 1013 | 304 | 2 |
Appendix B Examples
B.1 Success Cases
B.2 Failure Cases